Farnham Maintenance Reports

The dates in bold are the dates of the council meetings that received the reports.

4th September 2017  Council meeting

Management Work 2016/2017 as reported by June Atkinson is as follows:

March :   The front of the old section of hide was re-cladded using floor boarding, then preservative was applied.

April:  New wood was placed around the edges of the tern platform.  Plastic netting was erected round the Pyramidal Orchid site. The ditch alongside the path to the hide was cleaned out. Fungicide was applied to the rear of the hide.

May:  The hide was treated with two coats of ‘Fence Life’ .  The Iris bed in front of the Phragmites was treated with a mixture of cooking oil and ‘Roundup’, applied with a paint brush. Rosebay Willowherb in the Pyramidal Orchid site was treated with ‘Roundup’ using a weed wiper.

June: Hogweed in Area 2 was treated with ‘Roundup’ applied with a paintbrush.

July: Alders and Willowherb on the top car park were sprayed with ‘Roundup’ but the work was not completed. Ragwort growing in several areas was cut.  Paths on the west side were strimmed.  Spraying of Alders on the top car park continued.  Whilst applying wood preservative to the old timber on the rear of the new section of the hide, I removed a rotting floorboard at ground level to find that the telegraph pole which was the main rear support for the hide was rotting away because of rain falling from the roof. I asked a member of the Angling Club, who was an executive for McAlpine for advice. He kindly did the work on it for us by inserting long bolts around the pole then building concrete up the pole from the base.  He also re-bolted three benches by the hide and replaced rotten wood on one of the seats.  The Pyramidal Orchids in the fenced off area produced 40 flower spikes, with an additional six outside the area.

15th March 2016   Council meeting

7th February 2016    A fibreglass kit was used to repair a small hole in the boat.  Thanks to Mike Smithson and his contacts, a load of bark chippings was delivered and this was spread on the path to the copse and in front of the bird feeding area.

 14th February 2016  Strimming of the top car park, begun in January, was completed.  Little maintenance was required on the top path to the eastern boundary as two sessions were undertaken in the previous year, a few branches and brambles were cut back.  Previous bracken spraying has been successful and will be reviewed this year to ascertain if further treatment will be necessary.  The paths to and around Area 5 were checked, brambles and overhanging branches were cut back. Blackthorn from the copse is encroaching above areas 2 and 3, a habitat favoured by some ground nesting birds.  The blackthorn was reduced to keep the area open and more attractive.  Bark chippings were spread on the path below the copse. Nest boxes were inspected, cleaned out and six were replaced.  We are grateful to Brian Derbyshire who supplied 12 new boxes.  The west side of the hide was treated with wood preservative.

13th March 2016  Twelve bags of sand were taken to the island to top up the Sand Martin wall.  The vegetation on the top car park, previously strimmed, was raked up.  Several alder and sycamore saplings were cut down and treated.  Area 1 is undergoing a makeover.  Last year, areas of open soil were seeded with fine grasses which have grown successfully.  Another area is now being prepared for similar treatment.  A top straining wire was inserted in the rabbit-proof netting on Areas 3 and 6 to prevent it from falling over.  Some overhanging hawthorn branches near the nest boxes in the west field were cut back.

David Holmes has agreed to join the Management Committee from this year.

Management Committee

Chair – June Atkinson

Members – Robert Brown, David Holmes, Dr. John R. Mather (advisory), Mike Smithson, Richard Stobbs.

 

25th January 2016      Council meeting

22nd November 2015
All areas strimmed on the last work day were raked off. Overhanging willows behind the pond were cut back and two large willows in front of the pond were pollarded. Branches from a large willow and a blackthorn overhanging the path below the copse were cut back. The strimmer-mower was used in the top car park to cut down rosebay willowherb; this work will continue.
27th December 2015.
All brash from previous work days was burnt, which necessitated moving the rubbish out of the water onto dry land as the lake level had risen. Area 5 was raked off.
17th January 2016
Branches of willows which were overhanging the West Path were cut back to the main trunks. The hedge alongside the pond was cut and raked off. Tall hawthorn and elder trees at the feeding station were cut back. Molehills were levelled and rabbit holes filled in around the feeding station in order to prevent accidents.

17th November 2015

Summer Spraying Programme

June 2015

  1. The brambles and birches along the path at the bottom of the mound leading to the East Bank were treated. Brambles alongside Area 5 were also treated.
  2. Alder, willow and birch in the area around the car park pond were controlled.
  3. Spraying of Alchemilla mollis on the top car park was begun.

July2015

  1. The spraying of Alchemilla mollis was completed.

August 2015

  1. Alder and birch along the edges of the West Side path up to Sandy Bank were sprayed. Further control of alder saplings in the car park and shelter bank took place.

September 2015.

  1. During two sessions, a Weed Wiper was used in the birch area on the West Side. The small shed was re-roofed with felt by Barry Dawson.

Team Management Work

Sunday 20th September

Crack and white willows on the back spit island were cut down and the stumps treated. Three new fence posts were erected to support the netting extension around Area 6. Willows around Area 1 were cut back. Alder and birch saplings along the lake side below the hide were cut down and the stumps treated.

Sunday 25th October . Work on cutting back willows below the hide was completed. Areas 1,2,3 and 6 and round the pond were strimmed. Several tall willows next to Area 2 were topped back. A small team replaced two retaining boards on the steps below the hide.

30th March 2015      Council meeting

Work carried out at Farnham GP by Management Team 14th – 29th March 2015.
I have been concerned for some time about the amount of moss which is affecting most of the site, in particular around the pond where it is so dense that natural regeneration from seeds is not taking place. Together with Mike Smithson, I began a trial using a springbok rake to remove the moss and create areas of open, bare earth. We will monitor any development over the summer months.

Management day 29th March 2015
Two members spent the whole day boarding out a boat, which had been given to us by the Angling Club, to create another nesting site for Common Terns. The boat will be moored to the right of the tern platform through the summer and brought in when the terns have finished breeding.

  1. A bramble patch by the lakeside below the hide was reduced as it was encroaching on the access to the boat launching site.
  2. The bank above Areas 3 and 5 was cleared of brambles which were encroaching on the footpath, and debris from a dead hawthorn was raked out and removed.
  3. Ash and alder saplings on the top car park were cut down and treated and the tops removed.
  4. In Area 1, bramble shoots were dug up and problem weeds were removed. Also in the rabbit netting around Area 1 were repaired. The first phase of revegetating Area 1 was begun as it had become overgrown with unwanted species. A small area had been treated last year to eliminate rosebay willowherb, spear thistles were also removed, but the main species were retained. The area has been sown with a diverse meadow grass seed of 14 species including three fescue species and catstail, foxtail and cocksfoot. This seemed the best mixture to encourage a good diversity of insects and we will minitower the results carefully. Seedlings of knapweed and devil’s bit scabious from other parts of the site were transplanted into Area 1.
  5. An extension to increase the height of rabbit netting around Area 6 was put in place to prevent them from jumping into the fenced area from the adjacent bankside.
  6. A fallen, dead hawthorn behind the fence at the pond was cut up and removed.

This work took 10 people five hours to complete.

It was the last job of the eight planned management days, though there are still a few jobs to be completed.

June E. Atkinson      Chairman, Management Committee

23rd February 2015      Council meeting

Management Day 8th February 2015
Work continued around the car park pond – all small alder saplings were cut down and treated. This area must be kept open to maintain the path giving access to the lake, for transporting materials to the island which have been delivered to the car park.
The area was raked and all the rubbish burnt.
The hedge at the pond was trimmed.
Work was begun on the top car park area to control the brambles.
Old tree stumps from work carried out in previous years, were levelled to make it safer and easier to use the strimmer mower.
Six management days have now been completed.
On Thursday 19th February, the large unsafe willow tree in the copse was felled, leaving a stump of about seven feet in height.

20th January 2015       Council  meeting

i) 14th December 2014  Branches overhanging the West Side path were cut back, alder saplings and aspen suckers were cut down and treated.  Ditches along the main track and the path to the hide were cleaned out.  Alder saplings are becoming a major problem on site and work has begun around the car park pond which is fast becoming a forest; alders have been cut down and treated, the work will continue.
ii) 11th January 2015  Work around the car park continues. The top car park area was cut with the strimmer mower.  Rain stopped the work and it will continue in February.

11th November 2014

No meeting in July/August 2014.  The following is the recent update.  The photos are of recent clearance work on the west side of the lake.
Click on a picture to enlarge.  Use browser back button to return to here.

 

Work carried out at Farnham GP by Management Team -September, October, November 2014

Sunday 21st September

The path along the East Bank was cut.  The willows below the hide and below Area 1 were cut back.  Areas 2, 3 6 and around the pond were strimmed and raked off.

Sunday 19th October

The large willow FGP Work Party (3)below the copse which was overhanging the path was cut back.

FGP Work Party (2)

Area 5 was strimmed and raked off.  On the west side, willows were cut back from the path.

Several birch and alder saplings were thinned out from the lakeside tree belt.  The lakeside willows were cut back.

 

Tuesday 4th NovemberFGP Work Party

The task of applying wood preservative to the hide was begun.

 

 

 

June 30th 2014

Farnham Update .Management day 20th June. Due to the wet summers of the last two years, the spraying programme has been delayed and, as a result, many birch and alder seedlings had spread and the saplings increased in size. During June, I sprayed to control the many birch and alder saplings, also brambles, in the open areas on the west side. This is an ongoing problem but necessary to prevent these invasive species spreading into areas of botanical interest.  The East Bank was cut with a strimmer mower.  The two diseased elm trees were felled on 18th June and burnt where they fell. I have also sprayed bramble growth and alders on the top car park area which is always a productive site for butterflies.  During the next week, paths will be cut prior to Members’ Day. 44 holes are occupied in the Sand Martin wall; fledging is expected this week.  7 pairs of Common Terns were present; three of which reared six young, two pairs are still on eggs and two pairs failed.

May 19th 2014

Farnham Update.  Management Day 23rd March.  Sand was taken out to the Sand Martin wall to top up the shortfall. A new membrane and carpet was laid on top of it. Some brambles were scythed down. A second team worked on the top car park where brambles were lopped off prior to the area being strimmed.  Along the top path of the East Bank brambles, hawthorn and blackberries were cut back in order to keep the path open.              Management Day 6th April. An extra day was called to finish the work on the East Bank path, where a large area of brambles were strimmed well back. Brambles were strimmed from the paths around Area5. The banksides of the lakeside pool along the West Path were cleared. The hedge at the pond was trimmed and the ditch beside the path up to the hide was cleared out. An estimated 15 pairs of Sand Martins are presently visiting the wall and another pair of Common Terns makes a total of 6.

 

March 4th 2014

Farnham Update . Management Day in January was cancelled due to bad weather and re-scheduled for 2nd February.  All branches and debris from the December work day were removed and burnt.  In the old copse, all overhanging branches reaching to the ground were cut back.  All wind blown debris and fallen branches were raked up.  Management Day 23rd  February.  Around the pond and Areas 2,3 and 6 were strimmed and raked off. New bramble growth in the top car park was cut back prior to strimming.  Bramble control took place on the East Bank top path; this involved cutting the brambles well back from the path edge.  Due to a strong wind the covers could not be taken out to be placed on the tern platforms. Nest boxes were inspected and cleaned out.  Six new boxes for Tree Sparrows were erected; we are very grateful to Brian Darbyshire who made the boxes for Farnham to replace the original ones used in the survey, as they are falling apart.  On 24th February Barry Dawson and Robert Brown put in an extra day to take out the covers to the tern platforms and carry out essential repairs.

New Path March 2014

New Path March 2014 Sue Coldwell

On 1st March, a long awaited project to level a very badly sloping path below the copse was begun. JEA started by cutting back invasive blackthorn and raking it off before Barry Dawson started to dig out and level the path, using boards to form the edging.  Large quantities of wood chippings were delivered free, on site, last week. These will be used for the footpath.

 

Jan 20th 2014

Farnham Update.  On the Work Day on 27th October 2013 the cutting back of  willows below the hide, begun last month, was finished. Two large willows by the pond were cut back.  All rubbish was burnt on the bonfire.  On the Work Day on 24th November 2013 Willows to the left of the hide were cut back, also hawthorns from around the feeding station.  Area 5 was strimmed and raked off. The application of wood preservative to the hide was completed. In December,  David Chandler who is qualified and licensed to use a chainsaw, cut up the willow tree which had split and fallen in the copse and then removed it.  He also cut down a silver birch tree which was overhanging the west side path.

Sept 23rd  2013

Farnham Report.  There is no report this month as the Autumn planning meeting is still to occur. Ann requested that the Farnham Ecological Report is on the next agenda when she has had time to read it.

July 22nd 2013

Farnham Update.       Management work completed before Members’ Day on 14th July.       All the paths were strimmed; the East Bank paths to the hide, around Area 5 and the South Bank path.  Overhanging branches were cut back along bramble patches, giving access to view White -letter Hairstreaks.  Branches overhanging the path to Area 5 were cut back.  Round-up control for hogweed, bramble and willow was used in Areas 1,3 and 6. Thanks to good weather, about 30 members  attended Members’ Day this year.  An optics and camera display, supplied by Marcus Grover from Northallerton, provided additional interest for those seeking new equipment.  Butterfly species produced an excellent display with many Meadow Browns, Gatekeepers and Ringlets recorded.  The White Letter Hairstreaks obliged, giving good opportunity for photographers. Dragonfly species were also numerous with Emperor and Black Tailed Skimmer, Brown Hawker and four species of damselfly. Many thanks go to David Alred who helped members to identify each species. The Common Terns have young on the islands, which are beginning to appear again as the water level recedes. This report will go on the website ( see NEWS items – members day )

March 4th 2013

Farnham Update  The Management Day on 27th January was cancelled because of the snow.  On 24th February, one group removed encroaching saplings and cut back overhanging branches from the path on the East Bank, they also removed a large blackthorn bush.  The other group strimmed and raked off the fenced-off areas and around the pond.  June also informed Council about the very high level of the Great Crested Newt pond and the problem which the management team now faces when they attempt to remove large volumes of curly water thyme which is now covering the pond.  The newts need some open water and the chemical herbicide which has been used successfully in the past is no longer available and so June is hoping that the water level will drop and the margins of the pond will dry out in the summer.  They will then be able to remove the vegetation and the soil into which it is anchored so that the pond may be restored.  Council discussed the matter and Jill suggested investigating the Pondlife website for any advice they may give

January 14th 2013

 Farnham Update   Sunday 25th November Management  Day   Only 6 members of the Management Team were present, as the weather was poor with light rain. Work begun on the previous work day below the hide, cutting back the willow growth, was completed.  The top path on the East Bank was cut using a strimmer mower.  A short work day!


Supported by:
Change Font Size